# On the Capacity of Self-Attention

### v1

@@ B.2 Construction II: Permutations Under Compressive Embeddings
Theorem B.2 (Multi‑head recognition under Gaussian unit‑norm embeddings).
Assume the setup and construction of Algorithm 1 with $h=\frac{m}{d_{\mathrm{model}}}$ heads, per‑head dimension $d_{k}=C\log m$ for $C$ sufficiently large, and $\tau=\tfrac{1}{2}d_{k}$. If $d_{\mathrm{model}}\geq c_{0}\log m$ for a sufficiently large absolute constant $c_{0}$, then with probability at least $1-m^{-3}$ over the draw of $(X,W_{\mathrm{sig}})$,
$\forall i\in V\;\exists\,k\in[h]\text{ with }i\in V_{k}:\quad S^{(k)}_{i,\pi(i)}>\tau\quad\text{and}\quad S^{(k)}_{ij}<\tau\ \ \forall j\neq\pi(i).$
Consequently, max‑pooling over heads correctly recognizes all edges and $D_{K}=h\,d_{k}=\Theta\!\Big(\frac{m\log m}{d_{\mathrm{model}}}\Big).$
@@
Let $\mathbf{u}_{i}:=\mathbf{x}_{i}X^{\top}=\mathbf{e}_{i}(XX^{\top})$ and $\boldsymbol{\delta}_{i}:=\mathbf{u}_{i}-\mathbf{e}_{i}\in\mathbb{R}^{m}.$ Thus $\mathbf{u}_{i}$ is the ii‑th row of the Gram matrix $G:=XX^{\top}$; it satisfies $\mathbf{u}_{i}(i)=1$ and, for $j\neq i$, $\mathbf{u}_{i}(j)=\langle\mathbf{x}_{i},\mathbf{x}_{j}\rangle$. Fix a head $k$ and a source $i\in V_{k}$. As in Construction I, write
@@
$\quad+\underbrace{\Big(\sum_{s\in V_{k}}\delta_{i,s}\mathbf{w}_{\pi(s)}\Big)\cdot\mathbf{w}_{j}\cdot\mathbb{I}(j\in T_{k})}_{N_{2}}+\underbrace{\Big(\sum_{s\in V_{k}}\delta_{i,s}\mathbf{w}_{\pi(s)}\Big)\cdot\Big(\sum_{t\in T_{k}}\delta_{j,t}\mathbf{w}_{t}\Big)}_{N_{3}}.$
@@
For Item 3, define independent mean‑zero sub‑exponential variables $Y_{a}:=\langle\mathbf{x}_{i},\mathbf{x}_{\pi^{-1}(a)}\rangle\cdot\langle\mathbf{x}_{j},\mathbf{x}_{a}\rangle$ for $a\in T_{k}$. Each has $\psi_{1}$‑norm $\Theta(1/d_{\mathrm{model}})$ and $\mathbb{E}Y_{a}=0$. Bernstein’s inequality implies $\Pr\big[\big|\sum_{a\in T_{k}}Y_{a}\big|\geq t\big]\leq 2\exp(-\Omega(\min\{d_{\mathrm{model}}t^{2},\ d_{\mathrm{model}}t\})).$ Taking $t=C_{3}\sqrt{(\log m)/d_{\mathrm{model}}}$ and union bounding over all $i,j,k$ proves Item 3 for $c_{0}$ large enough. Item 1 is immediate from unit‑norm rows. ∎
@@
$C_{3}\,\sqrt{\tfrac{\log m}{d_{\mathrm{model}}}}\ \leq\ \tfrac{1}{16}\qquad\text{and}\qquad(C_{\star}+2C_{4}\sqrt{C_{2}}+C_{5}C_{2})\sqrt{\tfrac{\log m}{d_{k}}}\ \leq\ \tfrac{1}{16}.$
@@
$S^{(k)}_{i,\pi(i)}\;\geq\;d_{k}\ -\ \underbrace{(C_{4}\sqrt{C_{2}\,d_{k}\log m})}_{|N_{1}|}\ -\ \underbrace{(C_{4}\sqrt{C_{2}\,d_{k}\log m})}_{|N_{2}|}\ -\ \underbrace{\big(C_{5}C_{2}\sqrt{d_{k}\log m}+\tfrac{1}{16}d_{k}\big)}_{|N_{3}|}\;>\;\tfrac{3}{4}d_{k}\;>\;\tau.$
@@
$|S^{(k)}_{ij}|\ \leq\ C_{4}\sqrt{C_{2}\,d_{k}\log m}+\big(C_{5}C_{2}\sqrt{d_{k}\log m}+\tfrac{1}{16}d_{k}\big)\ <\ \tfrac{1}{4}d_{k}\ <\ \tau.$
@@
This bound is asymptotically optimal, matching our lower bound within a constant factor. In the proof of Theorem 4.2, the non-edge bounds hold uniformly over all $(i,j,k)$ (we union bound over $(i,j,k)$), so for any $j\neq\pi(i)$ we have $S_{ij}^{(k)}<\tau$ for all heads $k$ and hence $S_{ij}^{\max}<\tau$, while the target edge satisfies $S_{i,\pi(i)}^{\max}>\tau$. Lemma 4.1 then yields correctness on $E|_{\mathcal{C}}$ for every context $\mathcal{C}$.
@@ B.3 Construction III: More General Embeddings
$X_{\mathrm{inv}}:=\frac{1}{\mu}X^{\top},\qquad\mathbf{u}_{i}:=\mathbf{x}_{i}X_{\mathrm{inv}}=\frac{1}{\mu}\mathbf{e}_{i}(XX^{\top}),\qquad\boldsymbol{\delta}_{i}:=\mathbf{u}_{i}-\mathbf{e}_{i},$
@@
$W^{\prime}_{Q,(k)}(i,:)\;=\;\begin{cases}\mathbf{w}_{\pi(i)}&i\in V_{k}\\[2.0pt] 0&\text{else}\end{cases},\qquad W^{\prime}_{K,(k)}(j,:)\;=\;\begin{cases}\mathbf{w}_{j}&j\in T_{k}\\[2.0pt] 0&\text{else}\end{cases}.$
@@
$\forall i\in V\;\exists\,k\in[h]\text{ with }i\in V_{k}:\quad S^{(k)}_{i,\pi(i)}>\tau\quad\text{and}\quad S^{(k)}_{ij}<\tau\ \ \forall j\neq\pi(i).$
@@
$\mathbf{u}_{i}(i)\in[1-\varepsilon_{d},1+\varepsilon_{d}]\ \text{ with }\ \varepsilon_{d}\lesssim\frac{1}{\sqrt{\mu}},\qquad\rho\ \lesssim\ \frac{B}{d_{\text{model}}},\qquad\gamma\ \lesssim\ B\,p_{B}^{2}.$
@@
$\sum_{s\in S}\!\langle\mathbf{x}_{i},\mathbf{x}_{s}\rangle^{2}\ \leq\ \sum_{s\in S}\!\langle\mathbf{x}_{i},\mathbf{x}_{s}\rangle\quad\text{and}\quad\mathbb{E}\big[\langle\mathbf{x}_{i},\mathbf{x}_{s}\rangle\big]=d_{\text{model}}p_{B}^{2},$
@@ B.4 Construction IV: General Graphs
$\big(W^{\prime}_{Q,(k)}\big)_{i,\cdot}:=\begin{cases}\mathbf{w}_{\pi_{k}(i)}&i\in V_{k}\\ \mathbf{0}&\text{otherwise}\end{cases},\qquad\big(W^{\prime}_{K,(k)}\big)_{j,\cdot}:=\begin{cases}\mathbf{w}_{j}&j\in T_{k}\\ \mathbf{0}&\text{otherwise}\end{cases},$
@@
A union bound over all heads and all pairs in those heads costs only a $\log$ factor absorbed by $d_{k}=C\log m$: using $|M_{k}|\leq d_{\text{model}}$ and $\sum_{k}|M_{k}|=m^{\prime}$, we have $\sum_{k}|M_{k}|^{2}\leq d_{\text{model}}\sum_{k}|M_{k}|=d_{\text{model}}m^{\prime}=O(m^{\prime}d_{\text{model}})$ events in total. Finally, max pooling across heads preserves separation (non‑edges are below $\tau$ in every head, and each true edge belongs to exactly one $M_{k}$), and Lemma 4.1 yields context‑robustness for arbitrary subsets $\mathcal{C}\subseteq V$. ∎
@@
Appendix C Additional Justification for the Model (Section 3)
Computational footprint.
While it is natural to consider the number of heads ($h$) and the per-head key/query dimension ($d_{k}$) as two separate resources, we argue that the most relevant complexity measure is their product. In practice, the computation for multiple heads is not performed as $h$ distinct operations but as a single, larger batched operation. Let $\mathbf{X}\in\mathbb{R}^{\ell\times d_{\text{model}}}$ stack the context embeddings. With $W_{Q}^{\text{cat}},W_{K}^{\text{cat}}\in\mathbb{R}^{d_{\text{model}}\times(hd_{k})}$ formed by concatenating head weights, queries/keys are $\mathbf{Q}_{\text{total}}=\mathbf{X}W_{Q}^{\text{cat}}$ and $\mathbf{K}_{\text{total}}=\mathbf{X}W_{K}^{\text{cat}}$. Both flops and parameter/memory cost scale as $O(\ell\,d_{\text{model}}\,hd_{k})$ and $O(d_{\text{model}}\,hd_{k})$, respectively, motivating $D_{K}=hd_{k}$ as the budget.
While sub-cubic matrix multiplication algorithms could theoretically make one large head asymptotically faster than several smaller ones, this effect is absent in practice. The true bottleneck for these operations on modern hardware is memory bandwidth—the rate at which the matrices can be fetched from memory. The total data moved is proportional to the size of the weight matrices, which is in turn proportional to $d_{\text{model}}\cdot(h\cdot d_{k})$. Because deep learning libraries are highly optimized to perform these batched multiplications, the performance typically tracks the total size of the key and query matrices, regardless of the number of heads.
Analyzing the QK channel in isolation.
Since RGR asks where a source should connect, the OV pathway only reweights or propagates information after the routing decision has been made. Thus, a correct edge can only dominate if the QK gate already concentrates sufficient mass on the true target. Increasing value dimension $D_{V}$ amplifies whatever QK selects; it does not fix mis‑routing. [^5] Note that our model is scoped to a single self-attention layer; multi-layer iterative routing is outside our abstraction. Furthermore, our construction aligns with recent mechanistic analyses that separate each attention head into an OV circuit (what is read/written) and a QK circuit (where to attend); the QK circuit determines the attention pattern and thus the directed edges in RGR [33, 19].
Aggregating across heads.
This ‘OR‑of‑heads’ max is an analysis device for the edge test; it does not assert cross‑head QK interaction in standard MHA implementations (where heads are combined in the value pathway). However, aggregating multi-head QK scores by a max implements the intended RGR semantics: each head $k$ specializes to a relational template, and an (untyped) edge should exist if any template fires, i.e., $S^{\max}_{pq}=\max_{k}S^{(k)}_{pq}>\tau$ (an OR-of-relations). If a smooth alternative is desired, replacing $\max$ with log-sum-exp preserves the binary decision up to a global threshold shift, since for any scores $a_{1},\dots,a_{h}$,
$\max_{k}a_{k}\;\leq\;\operatorname{LSE}(a)=\log\!\sum_{k=1}^{h}e^{a_{k}}\;\leq\;\max_{k}a_{k}+\log h,$
so one can retune $\tau$ by at most $\log h$ without changing the classifier [6].
From a detection viewpoint, $\max$ is the standard OR-of-detectors aggregator (as in max-pooling): it passes through the strongest localized evidence while suppressing clutter, whereas averaging or summing allows many weak, unrelated heads to “conspire” to cross the threshold—undesirable for a yes/no edge test. This aligns with empirical observations that only a small number of specialized heads dominate while many can be pruned with little effect; a max aggregator naturally yields a principled winner-take-all over these specialists without paying for redundant heads.
Thresholding vs. Softmax.
Our model makes binary edge decisions from the raw QK scores $S_{ij}$. As a result, our thresholding decision paradigm can be replaced with the usual $1/\sqrt{d_{k}}$ scaling and softmax without changing the asymptotic budgets used by our constructions for all but very dense graphs. Throughout this paper, we keep the thresholding rule for simplicity and ease of exposition, but we describe how to incorporate this change into our results.
Scaling. If the decision is made directly on raw scores, the $1/\sqrt{d_{k}}$ factor can be absorbed into the threshold on $S_{ij}$. If the decision is made after softmax, we absorb $1/\sqrt{d_{k}}$ into the learned projections $W_{Q},W_{K}$ so that the effective logits are unchanged.
Softmax. Let $a_{ij}=\exp(S_{ij})/\sum_{t\in\mathcal{C}}\exp(S_{it})$ be the row-wise softmax over a context $\mathcal{C}$ (we assume distinct members of $V$ in any context, so $|\mathcal{C}|=\ell\leq m$). We use a single global (per graph) threshold $\hat{\tau}$ on the attention weight:
$\text{decide $(i,j)$ is an edge}\quad\Longleftrightarrow\quad a_{ij}\ \geq\ \hat{\tau}\quad\Longleftrightarrow\quad S_{ij}-\operatorname{lse}_{t\in\mathcal{C}}S_{it}\ \geq\ \log\hat{\tau}$.
Let $\Delta$ be the maximum out-degree of the graph, and let
$\gamma\ :=\ \min_{i}\Big(\min_{j\in N_{i}}S_{ij}\ -\ \max_{j\notin N_{i}}S_{ij}\Big)$
be the uniform score gap. A standard bound [6] yields, for any positive $j\in N_{i}$ and any $\ell\leq m$,
$a_{ij}\ \geq\ \frac{1}{\,\Delta+(\ell-\Delta)e^{-\gamma}\,}.$ (11)
Therefore a single global threshold $\hat{\tau}$ works for all contexts $\ell\leq m$ provided
$\gamma\ \geq\ \log\!\Big(\frac{m-\Delta}{\,1/\hat{\tau}-\Delta\,}\Big)\qquad\text{and necessarily}\qquad\hat{\tau}\ \leq\ \frac{1}{\Delta}.$ (12)
The condition $\hat{\tau}\leq 1/\Delta$ is unavoidable because the $k$ positive weights in a row sum to at most $1$. If $\Delta\leq m^{1-\varepsilon}$ for some constant $\varepsilon\in(0,1]$ and we take $\hat{\tau}=c/\Delta$ with any fixed $c\in(0,1)$, then
$\gamma\ \geq\ \log\!\Big(\frac{m-\Delta}{\,\Delta(1/c-1)\,}\Big)\ =\ \log\!\Big(\tfrac{m}{\Delta}-1\Big)\;+\;\log\!\Big(\tfrac{c}{1-c}\Big)\ \geq\ \varepsilon\log m\;+\;O(1).$
In all of our signature-based constructions, the gap satisfies $\gamma=\Theta(d_{k})$ with high probability; hence choosing $d_{k}=C\log m$ with $C$ large enough to meet (12) yields a valid single global threshold $\hat{\tau}$ for all contexts $\ell\leq m$.
We also note that sparse alternatives to softmax (sparsemax and $\alpha$-entmax) implement threshold-like attention rules that align directly with our edge test [43, 48, 11].

### v3

@@ B.2 Construction II: Permutations Under Compressive Embeddings
Theorem B.2 (Multi-head recognition under Gaussian unit-norm embeddings, max-over-heads).
Assume Gaussian unit-norm embeddings with $d_{\mathrm{model}}\geq c_{0}\log m$ for a sufficiently large absolute constant $c_{0}$. Let $h=\frac{m}{d_{\mathrm{model}}}$ heads, per-head dimension $d_{k}=C\log m$ for a sufficiently large absolute constant $C$, and threshold $\tau=\tfrac{1}{2}d_{k}$. Construct $\{(W_{Q}^{(k)},W_{K}^{(k)})\}_{k=1}^{h}$ as in Algorithm 1 and let $k(i)$ denote the unique head index such that $i\in V_{k(i)}$.
Then with probability at least $1-m^{-3}$ over the draw of $(X,W_{\mathrm{sig}})$, simultaneously for all $i\in V$:
$S^{(k(i))}_{i,\pi(i)}>\tau\qquad\text{and}\qquad\max_{k\in[h]}\max_{j\neq\pi(i)}S^{(k)}_{ij}<\tau.$
Consequently, $\forall j\neq\pi(i),S^{\max}_{i,\pi(i)}>\tau>S^{\max}_{ij},$ so max-over-heads recovers all edges, and the total key budget satisfies $D_{K}=h\,d_{k}=\Theta\!\Big(\frac{m\log m}{d_{\mathrm{model}}}\Big).$
@@
Let $\mathbf{u}_{i}:=\mathbf{x}_{i}X^{\top}=\mathbf{e}_{i}(XX^{\top})$ and $\boldsymbol{\delta}_{i}:=\mathbf{u}_{i}-\mathbf{e}_{i}\in\mathbb{R}^{m}.$ Thus $\mathbf{u}_{i}$ is the ii‑th row of the Gram matrix $G:=XX^{\top}$; it satisfies $\mathbf{u}_{i}(i)=1$ and, for $j\neq i$, $\mathbf{u}_{i}(j)=\langle\mathbf{x}_{i},\mathbf{x}_{j}\rangle$. Fix a head $k$ and a source $i\in V_{k}$. Analogously to Construction I, write
@@
$\quad+\underbrace{\Big(\sum_{s\in V_{k}}\delta_{i,s}\mathbf{w}_{\pi(s)}\Big)\cdot\mathbf{w}_{j}\cdot\mathbb{I}(j\in T_{k})}_{N_{2}}$
$+\underbrace{\Big(\sum_{s\in V_{k}}\delta_{i,s}\mathbf{w}_{\pi(s)}\Big)\cdot\Big(\sum_{t\in T_{k}}\delta_{j,t}\mathbf{w}_{t}\Big)}_{N_{3}}.$
@@
For Item 3, define independent mean‑zero sub‑exponential variables $Y_{a}:=\langle\mathbf{x}_{i},\mathbf{x}_{\pi^{-1}(a)}\rangle\cdot\langle\mathbf{x}_{j},\mathbf{x}_{a}\rangle$ for $a\in T_{k}$. Each has $\psi_{1}$‑norm $\Theta(1/d_{\mathrm{model}})$ and $\mathbb{E}[Y_{a}]=0$. Bernstein’s inequality implies $\Pr\big[\big|\sum_{a\in T_{k}}Y_{a}\big|\geq t\big]\leq 2\exp(-\Omega(\min\{d_{\mathrm{model}}t^{2},\ d_{\mathrm{model}}t\})).$ Taking $t=C_{3}\sqrt{(\log m)/d_{\mathrm{model}}}$ and union bounding over all $i,j,k$ proves Item 3 for $c_{0}$ large enough. Item 1 is immediate from unit‑norm rows. ∎
@@
$C_{3}\,\sqrt{\tfrac{\log m}{d_{\mathrm{model}}}}\ \leq\ \tfrac{1}{16}$
and
$(C_{\star}+2C_{4}\sqrt{C_{2}}+C_{5}C_{2})\sqrt{\tfrac{\log m}{d_{k}}}\ \leq\ \tfrac{1}{16}.$
@@
$S^{(k)}_{i,\pi(i)}\;\geq$ $d_{k}\ -\ \underbrace{(C_{4}\sqrt{C_{2}\,d_{k}\log m})}_{|N_{1}|}\ -\ \underbrace{(C_{4}\sqrt{C_{2}\,d_{k}\log m})}_{|N_{2}|}$
$-\ \underbrace{\big(C_{5}C_{2}\sqrt{d_{k}\log m}+\tfrac{1}{16}d_{k}\big)}_{|N_{3}|}\;>\;\tfrac{3}{4}d_{k}\;>\;\tau.$
@@
$|S^{(k)}_{ij}|\ \leq\ C_{4}\sqrt{C_{2}\,d_{k}\log m}+\big(C_{5}C_{2}\sqrt{d_{k}\log m}+\tfrac{1}{16}d_{k}\big)\$
$\tfrac{1}{4}d_{k}\ <\ \tau.$
@@
This bound is asymptotically optimal, matching our lower bound within a constant factor, since $m^{\prime}=m$ in the case of permutation graphs. In the proof of Theorem 4.1, the non-edge bounds hold uniformly over all $(i,j,k)$ (we union bound over $(i,j,k)$), so for any $j\neq\pi(i)$ we have $S_{ij}^{(k)}<\tau$ for all heads $k$ and hence $S_{ij}^{\max}<\tau$, while the target edge satisfies $S_{i,\pi(i)}^{\max}>\tau$. Monotonicity under context restriction then yields correctness on $E|_{\mathcal{C}}$ for every context $\mathcal{C}$.
@@ B.3 Construction III: More General Embeddings
$X_{\mathrm{inv}}:=\frac{1}{\mu}X^{\top},$
$\mathbf{u}_{i}:=\mathbf{x}_{i}X_{\mathrm{inv}}=\frac{1}{\mu}\mathbf{e}_{i}(XX^{\top}),$
$\boldsymbol{\delta}_{i}:=\mathbf{u}_{i}-\mathbf{e}_{i},$
@@
$W^{\prime}_{Q,(k)}(i,:)\;=\;\begin{cases}\mathbf{w}_{\pi(i)}&i\in V_{k}\\[2.0pt] 0&\text{else}\end{cases}$
$W^{\prime}_{K,(k)}(j,:)\;=\;\begin{cases}\mathbf{w}_{j}&j\in T_{k}\\[2.0pt] 0&\text{else}\end{cases}.$
@@
$\begin{split}&\forall i\in V\;\exists\,k\in[h]\text{ with }i\in V_{k}:\\ &\quad S^{(k)}_{i,\pi(i)}>\tau\quad\text{and}\quad S^{(k)}_{ij}<\tau\ \ \forall j\neq\pi(i).\end{split}$
@@
$\mathbf{u}_{i}(i)\in[1-\varepsilon_{d},1+\varepsilon_{d}]\ \text{ with }\ \varepsilon_{d}\lesssim\frac{1}{\sqrt{\mu}},$
$\rho\ \lesssim\ \frac{B}{d_{\text{model}}},$
$\gamma\ \lesssim\ B\,p_{B}^{2}.$
@@
$\sum_{s\in S}\!\langle\mathbf{x}_{i},\mathbf{x}_{s}\rangle^{2}\ \leq\ \sum_{s\in S}\!\langle\mathbf{x}_{i},\mathbf{x}_{s}\rangle$
and
$\mathbb{E}\big[\langle\mathbf{x}_{i},\mathbf{x}_{s}\rangle\big]=d_{\text{model}}p_{B}^{2},$
@@ B.4 Construction IV: General Graphs
$\big(W^{\prime}_{Q,(k)}\big)_{i,\cdot}:=\begin{cases}\mathbf{w}_{\pi_{k}(i)}&i\in V_{k}\\ \mathbf{0}&\text{otherwise}\end{cases},$
$\big(W^{\prime}_{K,(k)}\big)_{j,\cdot}:=\begin{cases}\mathbf{w}_{j}&j\in T_{k}\\ \mathbf{0}&\text{otherwise}\end{cases},$
@@
A union bound over all heads and all pairs in those heads costs only a $\log$ factor absorbed by $d_{k}=C\log m$: using $|M_{k}|\leq d_{\text{model}}$ and $\sum_{k}|M_{k}|=m^{\prime}$, we have $\sum_{k}|M_{k}|^{2}\leq d_{\text{model}}\sum_{k}|M_{k}|=d_{\text{model}}m^{\prime}=O(m^{\prime}d_{\text{model}})$ events in total. Finally, max pooling across heads preserves separation (non‑edges are below $\tau$ in every head, and each true edge belongs to exactly one $M_{k}$), and monotonicity under context restriction yields context‑robustness for arbitrary subsets $\mathcal{C}\subseteq V$. ∎
@@
Appendix C Additional Justification: Computational Footprint of the Model (Section 3)
While it is natural to consider the number of heads ($h$) and the per-head key/query dimension ($d_{k}$) as two separate resources, we argue that the most relevant complexity measure is their product
$D_{K}\;:=\;hd_{k},$
since standard implementations perform multi-head attention using batched dense linear algebra whose leading costs depend on $D_{K}$.
Concretely, let $\mathbf{X}\in\mathbb{R}^{\ell\times d_{\text{model}}}$ stack the context embeddings. With
$W_{Q}^{\text{cat}}=[W_{Q}^{(1)}|\cdots|W_{Q}^{(h)}]\in\mathbb{R}^{d_{\text{model}}\times(hd_{k})}$
and
$W_{K}^{\text{cat}}=[W_{K}^{(1)}|\cdots|W_{K}^{(h)}]\in\mathbb{R}^{d_{\text{model}}\times(hd_{k})}$
formed by concatenating the head weights, the total queries/keys are computed as
$\mathbf{Q}_{\text{total}}=\mathbf{X}W_{Q}^{\text{cat}}\qquad\text{and}\qquad\mathbf{K}_{\text{total}}=\mathbf{X}W_{K}^{\text{cat}}.$
Thus, both flops and parameter/memory cost for forming $\mathbf{Q}_{\text{total}}$ and $\mathbf{K}_{\text{total}}$ scale as
$O(\ell\,d_{\text{model}}\,hd_{k})=O(\ell\,d_{\text{model}}\,D_{K})$
and
$O(d_{\text{model}}\,hd_{k})=O(d_{\text{model}}\,D_{K}),$
respectively, motivating $D_{K}$ as the appropriate budget.
This same dependence on $D_{K}$ persists in the subsequent computation of per-head logits. Reshape $\mathbf{Q}_{\text{total}},\mathbf{K}_{\text{total}}$ into $\mathbf{Q}\in\mathbb{R}^{h\times\ell\times d_{k}}$ and $\mathbf{K}\in\mathbb{R}^{h\times\ell\times d_{k}}$, and form per-head score matrices
$\mathbf{S}^{(t)}\;=\;\mathbf{Q}^{(t)}\big(\mathbf{K}^{(t)}\big)^{\top}\in\mathbb{R}^{\ell\times\ell},\qquad t\in[h],$
(with the usual $d_{k}^{-1/2}$ scaling in the softmax variant, which does not change asymptotic cost). The total work to compute all $\{\mathbf{S}^{(t)}\}_{t=1}^{h}$ scales as
$O(h\,\ell^{2}\,d_{k})\;=\;O(\ell^{2}\,D_{K}).$
While sub-cubic matrix multiplication algorithms could theoretically make one large head asymptotically faster than several smaller ones, this effect is absent in practice. The dominant kernels here are highly optimized batched matrix multiplications (and, when applicable, softmax/normalization), and their observed throughput typically tracks the total matrix sizes involved. Consequently, at fixed $D_{K}=hd_{k}$, varying the split between $h$ and $d_{k}$ has little effect on the leading computational footprint of the $QK$ channel.